OpenStudy (anonymous):

The Frostburg-Truth bus travels on a straight road from Frostburg Mall to Sojourner Truth Park. The mall is 3 miles west and 2 miles south of the City Center. The park is 4 miles east and 5 miles north of the Center. How far is it from the mall to the park to the nearest tenth of a mile?

OpenStudy (anonymous):

Find the distance from (-5,-3) to (4,4) Distance= Sqrt[(-5-4)^2+(-3-4)^2]=Sqrt(81+49)= Sqrt(130) Without a calculator handy, the square root of 130 will be more than 11 (because 11*11=121) and less than 12 (12*12=144) but a little closer to 11 D. 11.4 miles (on a calculator 11.40175425099138)